# Inverse Hamiltonian reduction in type A and generalized slices in the affine Grassmannian

### v1

@@ 4.2. Strongly equivariant quantizations
An $H$-equivariant quantization $(A,\varphi)$ is said to be strongly $H$-equivariant if it has a comoment map, i.e., a $H$-equivariant morphism of associative $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$-algebras, $\Phi^{\#}:U_{\hbar}(\mathfrak{h})\rightarrow A$, such that for any $x\in\mathfrak{h}$, its image under the differential of the action, $d\rho(x)$ agrees with the derivation $[\Phi^{\#}(x),\cdot]$.
@@
We say that a map, ${\mathcal{A}}\rightarrow{\mathcal{A}}^{\prime}$, of strongly equivariant algebras (Poisson or associative) is strongly $H$-equivariant if it is $H$-equivariant and intertwines the comoment maps, i.e.the following diagram commutes
[…]
Such strongly equivariant quantizations have been classified, in the $C^{\infty}$ setting via Fedosov-style arguments by [RW16]. Recently, a deformation theory argument has appeared in [ENST25]
@@
The set of isomorphism classes of strongly equivariant graded quantizations is in bijection with the second $H$-equivariant cohomology
(4.2.2) $H^{2}_{H}(Y)~{},$
that is, any two quantizations $A$ and $A^{\prime}$ corresponding to the same equivariant cohomology class are isomorphic and the isomorphism is strongly $H$-equivariant.
@@ 4.3. Quantum Hamiltonian reduction
It is not true in general that $A_{\rm KW}$ is a quantization of $Y_{\rm KW}$ but for a nice enough $H$-action this is the case—we shall give sufficient conditions shortly.
@@
[…]
@@ 5. Inverse Hamiltonian reduction
Now let us specialize so that $H$ is abelian. The $\mathbb{\mathbb{C}}$-scheme $Y$ continues to be affine, symplectic, and equipped with a Hamiltonian action of $H$, with moment map $\Phi:Y\rightarrow\mathfrak{h}^{*}$. Moreover, we continue to assume that $Y$ is equipped with an action of $\mathbb{G}_{\hbar}$, that scales the symplectic form with weight $-1$.
@@
[…]
@@ 6. Multiplication of slices
To show that the equivariant Slodowy slices admit inverse Hamiltonian reductions, we shall make use of their incarnation as generalized slices in the affine Grassmannian. In the following section we introduce the necessary machinery to systematically construct inverse Hamiltonian reduction data for generalized slices in type $A$.
@@ 6.1. Partial convolution maps
For $\pi_{0}:\{1,...,n\}{\ \twoheadrightarrow\ }\{1,...,m\}$ and $\pi_{1}:\{1,...,m\}\to\{1,...,\ell\}$ order preserving surjections as above, the diagram
@@
The restriction along the diagonal embedding $\Delta(\pi):\mathbb{A}^{m}\to\mathbb{A}^{n}$ implies that $\sigma_{i_{j-1}+1}\circ\ldots\circ\sigma_{i_{j}}$ determines an isomorphism of principal $G$-bundles away from the single point $x_{j}$ for $j=2,...,m$, and determines a trivialization away from the single point $x_{1}$ for $j=1$. ∎
@@
For $\pi_{0}:\{1,...,n\}{\ \twoheadrightarrow\ }\{1,...,m\}$ and $\pi_{1}:\{1,...,m\}\to\{1,...,\ell\}$ order preserving surjections as above, the diagram
@@ 6.2. Symmetric convolution slices
  • $\mathcal{P}_{i}\in\operatorname{Bun}_{G}(\mathbb{P}^{1})$ is a principal $G$-bundle on $\mathbb{P}^{1}$ for $i=0,...,n$;
  • $\sigma_{i}:\mathcal{P}_{i}|_{\mathbb{A}^{1}_{\infty}}\xrightarrow{\cong}\mathcal{P}_{i-1}|_{\mathbb{A}^{1}_{\infty}}$ is an isomorphism of principal $G$-bundles with pole at 00 bounded by $\lambda_{i}$ for $i=1,...,n$;
@@
  • $\mathcal{P}_{B_{-}}\in\operatorname{Bun}_{B_{-}}^{-w_{0}\mu_{-}}(\mathbb{P}^{1})$ is a principal $B_{-}$-bundle on $\mathbb{P}^{1}$ of degree $-w_{0}\mu_{-}$; and
@@
[…]
@@
defined by modifying the underlying $G$-bundles by a Hecke transformation at $\infty$ of weight $-w_{0}\mu_{-}$.
@@ 6.3. Multiplication maps
as the composition ${\mathcal{W}}^{\lambda_{1}}_{\mu_{1}}\times{\mathcal{W}}^{{\underline{\lambda}}^{2}}_{\mu_{2}}\xrightarrow{\widetilde{\bf m}}\widetilde{{\mathcal{W}}}^{(\lambda_{1},{\underline{\lambda}}^{2})}_{\mu_{1}+\mu_{2}}\xrightarrow{\varpi^{\pi}}{\mathcal{W}}^{\lambda_{1}+\lambda_{2}}_{\mu_{1}+\mu_{2}}$, where $\varpi^{\pi}$ denotes the partial convolution map determined by the surjection $\pi:\{1,...,n\}\to\{1,...,n-1\}$ defined by $\pi(1)=1$ and $p(i)=i-1$ for $i\geq 2$.
@@ 6.4. Multiplication maps on Beilinson–Drinfeld generalized slices
  • $\mathcal{P}_{i}\in\operatorname{Bun}_{G}(\mathbb{P}^{1})$ is a principal $G$-bundle on $\mathbb{P}^{1}$ for $i=0,...,n$;
  • $\sigma_{i}:\mathcal{P}_{i}|_{\mathbb{P}^{1}\setminus\{x_{i}\}}\xrightarrow{\cong}\mathcal{P}_{i-1}|_{\mathbb{P}^{1}\setminus\{x_{i}\}}$ is an isomorphism of principal $G$-bundles with pole at $x_{i}$ bounded by $\lambda_{i}$ for $i=1,...,n$;
@@
  • $\mathcal{P}_{B_{-}}\in\operatorname{Bun}_{B_{-}}^{-w_{0}\mu_{-}}(\mathbb{P}^{1})$ is a principal $B_{-}$-bundle on $\mathbb{P}^{1}$ of degree $-w_{0}\mu_{-}$;
@@
[…]
@@
$\underline{\mathcal{W}}^{\lambda_{1}}_{\mu_{1}}\times_{\mathbb{A}^{1}}\underline{\mathcal{W}}^{{\underline{\lambda}}^{2}}_{\mu_{2}}\xrightarrow{\widetilde{\bf m}}\widetilde{\underline{\mathcal{W}}}^{(\lambda_{1},{\underline{\lambda}}^{2})}_{\mu_{1}+\mu_{2}}\times_{\mathbb{A}^{2}}\mathbb{A}^{1}\cong\widetilde{\underline{\mathcal{W}}}^{(\lambda_{1},{\underline{\lambda}}^{2})}_{\mu_{1}+\mu_{2}}\times_{\mathbb{A}^{n+1}}\mathbb{A}^{n}\xrightarrow{\varpi^{\pi}}\underline{\mathcal{W}}^{(\lambda_{1}+{\underline{\lambda}}^{2})}_{\mu_{1}+\mu_{2}}\ ,$
where $\varpi^{\pi}$ denotes the partial convolution map determined by the surjection $\pi:\{1,...,n\}\to\{1,...,n-1\}$ defined by $\pi(1)=1$ and $p(i)=i-1$ for $i\geq 2$.
@@ 7. Inverse Hamiltonian reduction for generalized slices
where $N_{1,k}[\![t^{-1}]\!]\subset N_{k}[\![t^{-1}]\!]$ is the first congruent subgroup of the upper unipotent subgroup of ${\rm GL}_{k}[\![t^{-1}]\!]$
@@ 7.1. An action of the translation group
Let $\mathfrak{G}_{\alpha}\coloneqq{\rm Lie}\,\mathbb{G}_{\alpha}$ thought of as a commutative, Lie subalgebra of $\mathfrak{n}$ generated by the roots $(\alpha_{r_{1}},\alpha_{r_{1}}+\alpha_{r_{1}+1},\dots,\sum_{i=r_{1}}^{r_{2}}\alpha_{i})$. Define $\psi_{\alpha}:N\rightarrow\mathfrak{G}_{\alpha}$ by
@@
This lifts, naturally to a map on the loop spaces $\psi_{\alpha}:N(\!(t^{-1)})\!)\rightarrow\mathfrak{G}_{\alpha}(\!(t^{-1})\!)$. Moreover, we can extend $\psi_{\alpha}$ to a map $\psi_{\alpha}:\mathrm{Gr}_{\mu}\rightarrow\mathfrak{G}_{\alpha}{{(\!(t)\!)}}$ by
@@
In particular, we use $\Phi_{\alpha}:N(\!(t^{-1})\!)\rightarrow\mathfrak{G}_{\alpha}$ for the restriction to the first Fourier modes, i.e.,
@@ 7.2. Action on Zastava
Suppose, $m<n-1$, then $\underline{y}_{n-m,\bullet}=(0,0,\dots,b_{m-1})$ and $\underline{y}_{\bullet,n-m}=(0,0,\dots,g_{m-1})^{\rm T}$. So from (7.2.6), we see that
@@
From 7.3 we see that the action of $\mathbb{G}_{\alpha}$ preserves the symplectic form and that the moment map is the projection to the $(b_{i},e)$ co-ordinates. What remains is to show that the map $\Phi_{\alpha}$ agrees with this projection but this is precisely the content of Lemma 7.4. ∎
@@
(7.2.17) $\overset{\circ}{\mathfrak{G}_{\alpha}}\equiv\Phi_{\alpha}(\mathrm{Gr}^{0}_{-\alpha})\subset\mathfrak{G}_{\alpha}~{}.$
@@
For convenience we write, $(\overset{\circ}{\mathfrak{G}}_{\alpha})_{\mu}$ for the preimage of $\overset{\circ}{\mathfrak{G}}_{\alpha}$ under $\Phi_{\alpha}:\mathrm{Gr}_{\mu}\rightarrow\mathfrak{G}_{\alpha}$.
Define a map $\xi_{\alpha}:(\overset{\circ}{\mathfrak{G}}_{\alpha})_{\mu}\rightarrow\mathrm{Gr}_{-\alpha}^{0}$ on the usual co-ordinates by
@@
Note that in $(\overset{\circ}{\mathfrak{G}}_{\alpha})_{\mu}$, $\Phi_{\alpha,k}(y)$ is nonzero, hence this map is well-defined.
@@
(7.2.19) $(\overset{\circ}{\mathfrak{G}_{\alpha}})_{-\alpha}^{0}\xrightarrow{\xi_{\alpha}}\mathrm{Gr}^{0}_{\alpha}$
@@
Proposition 7.5 implies that $(\overset{\circ}{\mathfrak{G}}_{\alpha})_{-\alpha}^{0}=\mathrm{Gr}_{-\alpha}^{0}$ and $\xi_{\alpha}(y)=y$ follows from Lemma 7.4 and the definition of $\xi_{\alpha}$. ∎

### v2

@@ 4.2. Strongly equivariant quantizations
An $H$-equivariant quantization $(A,\varphi)$ is said to be strongly $H$-equivariant if it has a comoment map, i.e., a $H$-equivariant morphism of associative $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$-algebras, $\Phi^{\#}:U_{\hbar}(\mathfrak{h})\rightarrow A$, such that for any $x\in\mathfrak{h}$, its image under the differential of the action, $d\rho(x)$ agrees with the derivation $\Phi^{\#}(x),\cdot]$.
@@
We say that a map, $A\rightarrow A^{\prime}$, of strongly equivariant algebras (Poisson or associative) is strongly $H$-equivariant if it is $H$-equivariant and intertwines the comoment maps, i.e.the following diagram commutes
(4.2.1) $A{A^{\prime}}$$U_{\hbar}(\mathfrak{h})$$U_{\hbar}(\mathfrak{h})$$\scriptstyle{\rm id}$
Such strongly equivariant quantizations have been classified in the $C^{\infty}$ setting via Fedosov-style arguments by [RW16]. Recently, a deformation theory argument has appeared in [ENST25]. However, the assumptions required in the $C^{\infty}$ setting do not survive into the algebraic setting [^2] The authors wish to thank I. Losev for bringing this to our attention
To this end, we prove a version of equivariant deformation theory in [BN25b], and we recall the main result here.
@@
Suppose a strongly $H$-equivariant, graded quantization of $Y$ exists. Then the moduli space of such quantizations, up to isomorphism, is a torsor for the second $H$-equivariant cohomology
(4.2.2) $H^{2}_{H}(Y)~.$
In particular, there exists a strongly $H$-equivariant, graded isomorphism between any two strongly $H$-equivariant, graded quantizations $A$ and $A^{\prime}$ with the same equivariant cohomology class.
Proposition 4.6 ([BN25b]).
Suppose $H$ is unipotent and abelian, and the Hamiltonian action of $H$ on $Y$ is such that
  1. (i)
    the action of $H$ is free, and
  2. (ii)
    the moment map $\Phi:Y\rightarrow\mathfrak{h}^{*}$ is smooth.
Then every graded quantization of $Y$ can be made strongly $H$-equivariant. Moreover, such a lift is unique up to strongly $H$-equivariant isomorphism.
@@ 4.3. Quantum Hamiltonian reduction
It is not true in general that $A/\!\!/\!\!/_{\chi}H$ is a quantization of $Y/\!\!/\!\!/_{\chi}H$. However, for a nice enough $H$-action this is the case—we shall give sufficient conditions shortly.
@@
Suppose $H$ acts on $Y$ freely and the moment map $\Phi:Y\rightarrow\mathfrak{h}^{*}$ is flat in a neighborhood around $\chi$. Then, if $({\mathcal{A}},\varphi,\Phi^{\#})$ is any strongly $H$-equivariant quantization of ${\mathcal{O}}_{Y}$, ${\mathcal{A}}\underset{\chi}{/\!\!/\!\!/}H$ is a quantization of ${\mathcal{O}}_{Y_{\rm KW}}$. Moreover, if $A$ is any strongly $H$-equivariant quantization of ${\mathcal{O}}(Y)$, then $A/\!\!/\!\!/_{\chi}H$ is a quantization of ${\mathcal{O}}(Y/\!\!/\!\!/_{\chi}H)$.
@@ 5. Inverse Hamiltonian reduction
Now let us specialize so that $H$ is abelian. The $\mathbb{\mathbb{C}}$-scheme $Y$ continues to be affine, symplectic, and equipped with a Hamiltonian action of $H$, with moment map $\Phi:Y\rightarrow\mathfrak{h}^{*}$. Moreover, we continue to assume that $Y$ is equipped with an action of $\mathbb{G}_{\hbar}$, that scales the symplectic form with weight 1-1.
@@
The pullback of functions along $H\times U\times Y/\!\!/\!\!/_{\chi}H\xrightarrow{f}Y\underset{\mu,\mathfrak{h}^{*}}{\times}U\rightarrow Y$ gives a strongly $H$-equivariant embedding ${\mathcal{O}}(Y)\hookrightarrow{\mathcal{O}}(H\times U)\otimes{\mathcal{O}}(Y/\!\!/\!\!/_{\chi}H)$ of Poisson algebras. Hence, the nomenclature inverse Hamiltonian reduction.
@@ 6. Multiplication of slices
To show that the equivariant Slodowy slices admit inverse Hamiltonian reductions, we shall make use of their incarnation as generalized slices in the affine Grassmannian. In the following section we introduce the necessary machinery to systematically construct inverse Hamiltonian reduction data for generalized slices in type $A$. A number of the results of this section on multiplication maps between generalized slices closely follow those from [BFN19a], [KP21].
@@ 6.1. Partial convolution maps
For $\pi_{0}:\{1,...,n\}{\ \twoheadrightarrow\ }\{1,...,m\}$ and $\pi_{1}:\{1,...,m\}\to\{1,$$\ell\}$ order preserving surjections as above, the diagram
@@
The restriction along the diagonal embedding $\Delta(\pi):\mathbb{A}^{m}\to\mathbb{A}^{n}$ implies that $\sigma_{i_{j-1}+1}\circ\ldots\circ\sigma_{i_{j}}$ determines an isomorphism of principal $G$-bundles away from the single point $x_{j}$ for j=2,,mj=2,...,m, and determines a trivialization away from the single point $x_{1}$ for $j=1$. ∎
@@
For $\pi_{0}:\{1,...,n\}{\ \twoheadrightarrow\ }\{1,...,m\}$ and $\pi_{1}:\{1,...,m\}\to\{1,$$\ell\}$ order preserving surjections as above, the diagram
@@ 6.2. Symmetric convolution slices
  • $\mathcal{P}_{i}\in\operatorname{Bun}_{G}(\mathbb{P}^{1})$ is a principal $G$-bundle on $\mathbb{P}^{1}$ for i=0,,ni=0,...,n;
  • $\sigma_{i}:\mathcal{P}_{i}|_{\mathbb{A}^{1}_{\infty}}\xrightarrow{\cong}\mathcal{P}_{i-1}|_{\mathbb{A}^{1}_{\infty}}$ is an isomorphism of principal $G$-bundles with pole at 0 bounded by $\lambda_{i}$ for i=1,,ni=1,...,n;
@@
  • $\mathcal{P}_{B_{-}}\in\operatorname{Bun}_{B_{-}}^{-w_{0}\mu_{-}}(\mathbb{P}^{1})$ is a principal $B_{-}$-bundle on $\mathbb{P}^{1}$ of degree $w_{0}\mu_{-}$; and
@@
    $\mathcal{P}_{B_{-}}|_{\infty}\to\mathcal{P}_{B_{-}}|_{\infty}\times_{B_{-}}G\xrightarrow{\phi_{-}|_{\infty}}\mathcal{P}_{0}|_{\infty}\xrightarrow{s}G$
@@
defined by modifying the underlying $G$-bundles by a Hecke transformation at $\infty$ of weight $w_{0}\mu_{-}$.
@@ 6.3. Multiplication maps
as the composition ${\mathcal{W}}^{\lambda_{1}}_{\mu_{1}}\times{\mathcal{W}}^{{\underline{\lambda}}^{2}}_{\mu_{2}}\xrightarrow{\widetilde{\bf m}}\widetilde{{\mathcal{W}}}^{(\lambda_{1},{\underline{\lambda}}^{2})}_{\mu_{1}+\mu_{2}}\xrightarrow{\varpi^{\pi}}{\mathcal{W}}^{\lambda_{1}+\lambda_{2}}_{\mu_{1}+\mu_{2}}$, where $\varpi^{\pi}$ denotes the partial convolution map determined by the surjection $\pi:\{1,...,n\}\to\{1,...,n-1\}$ defined by $\pi(1)=1$ and p(i)=i1p(i)=i-1 for $i\geq 2$.
@@ 6.4. Multiplication maps on Beilinson–Drinfeld generalized slices
  • $\mathcal{P}_{i}\in\operatorname{Bun}_{G}(\mathbb{P}^{1})$ is a principal $G$-bundle on $\mathbb{P}^{1}$ for i=0,,ni=0,...,n;
  • $\sigma_{i}:\mathcal{P}_{i}|_{\mathbb{P}^{1}\setminus\{x_{i}\}}\xrightarrow{\cong}\mathcal{P}_{i-1}|_{\mathbb{P}^{1}\setminus\{x_{i}\}}$ is an isomorphism of principal $G$-bundles with pole at $x_{i}$ bounded by $\lambda_{i}$ for i=1,,ni=1,...,n;
@@
  • $\mathcal{P}_{B_{-}}\in\operatorname{Bun}_{B_{-}}^{-w_{0}\mu_{-}}(\mathbb{P}^{1})$ is a principal $B_{-}$-bundle on $\mathbb{P}^{1}$ of degree $w_{0}\mu_{-}$;
@@
    $\mathcal{P}_{B_{-}}|_{\infty}\to\mathcal{P}_{B_{-}}|_{\infty}\times_{B_{-}}G\xrightarrow{\phi_{-}|_{\infty}}\mathcal{P}_{0}|_{\infty}\xrightarrow{s}G$
@@
$\underline{\mathcal{W}}^{\lambda_{1}}_{\mu_{1}}\times_{\mathbb{A}^{1}}\underline{\mathcal{W}}^{{\underline{\lambda}}^{2}}_{\mu_{2}}\xrightarrow{\widetilde{\bf m}}\widetilde{\underline{\mathcal{W}}}^{(\lambda_{1},{\underline{\lambda}}^{2})}_{\mu_{1}+\mu_{2}}\times_{\mathbb{A}^{2}}\mathbb{A}^{1}\cong\widetilde{\underline{\mathcal{W}}}^{(\lambda_{1},{\underline{\lambda}}^{2})}_{\mu_{1}+\mu_{2}}\times_{\mathbb{A}^{n+1}}\mathbb{A}^{n}\xrightarrow{\varpi^{\pi}}\underline{\mathcal{W}}^{(\lambda_{1}+{\underline{\lambda}}^{2})}_{\mu_{1}+\mu_{2}}\$,
where $\varpi^{\pi}$ denotes the partial convolution map determined by the surjection $\pi:\{1,...,n\}\to\{1,...,n-1\}$ defined by $\pi(1)=1$ and p(i)=i1p(i)=i-1 for $i\geq 2$.
@@ 7. Inverse Hamiltonian reduction for generalized slices
where $N_{1,k}[\![t^{-1}]\!]\subset N_{k}[\![t^{-1}]\!]$ is the first congruent subgroup of the upper unipotent subgroup of ${\rm GL}_{k}[\![t^{-1}]\!]$ .
@@ 7.1. An action of the translation group
Let $\mathfrak{G}_{\alpha}\coloneqq{\rm Lie}\,\mathbb{G}_{\alpha}$ thought of as a commutative, Lie subalgebra of $\mathfrak{n}$ generated by the roots $(\alpha_{r_{1}},\alpha_{r_{1}}+\alpha_{r_{1}+1},\dots,\sum_{i=r_{1}}^{r_{2}}\alpha_{i})$. Define $\psi_{\alpha}:N\rightarrow\mathfrak{G}_{\alpha}^{*}$ by
@@
This lifts, naturally to a map on the loop spaces $\psi_{\alpha}:N(\!(t^{-1)})\!)\rightarrow\mathfrak{G}_{\alpha}^{*}(\!(t^{-1})\!)$. Moreover, we can extend $\psi_{\alpha}$ to a map $\psi_{\alpha}:\mathrm{Gr}_{\mu}\rightarrow\mathfrak{G}_{\alpha}^{*}{{(\!(t)\!)}}$ by
@@
In particular, we use $\Phi_{\alpha}:N(\!(t^{-1})\!)\rightarrow\mathfrak{G}^{*}_{\alpha}$ for the restriction to the first Fourier modes, i.e.,
@@ 7.2. Action on Zastava
Suppose, m<n1m<n-1, then $\underline{y}_{n-m,\bullet}=(0,0,\dots,b_{m-1})$ and $\underline{y}_{\bullet,n-m}=(0,0,\dots,g_{m-1})^{\rm T}$. So from (7.2.6), we see that
@@
From 7.4 we see that the action of $\mathbb{G}_{\alpha}$ preserves the symplectic form and that the moment map is the projection to the $(b_{i},e)$ co-ordinates. What remains is to show that the map $\Phi_{\alpha}$ agrees with this projection but this is precisely the content of Lemma 7.5. ∎
@@
(7.2.17) $\overset{\circ}{\mathfrak{G}^{*}_{\alpha}}\equiv\Phi_{\alpha}(\mathrm{Gr}^{0}_{-\alpha})\subset\mathfrak{G}^{*}_{\alpha}~.$
@@
For convenience we write, $(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha})_{\mu}$ for the preimage of $\overset{\circ}{\mathfrak{G}^{*}}_{\alpha}$ under $\Phi_{\alpha}:\mathrm{Gr}_{\mu}\rightarrow\mathfrak{G}_{\alpha}$.
Define a map $\xi_{\alpha}:(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha})_{\mu}\rightarrow\mathrm{Gr}_{-\alpha}^{0}$ on the usual co-ordinates by
@@
Note that in $(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha})_{\mu}$, $\Phi_{\alpha,k}(y)$ is nonzero, hence this map is well-defined.
@@
(7.2.19) $(\overset{\circ}{\mathfrak{G}^{*}_{\alpha}})_{-\alpha}^{0}\xrightarrow{\xi_{\alpha}}\mathrm{Gr}^{0}_{\alpha}$
@@
Proposition 7.6 implies that $(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha})_{-\alpha}^{0}=\mathrm{Gr}_{-\alpha}^{0}$ and $\xi_{\alpha}(y)=y$ follows from Lemma 7.5 and the definition of $\xi_{\alpha}$. ∎